# Yoga for Fourier–Mukai partnership

### v2

@@ Acknowledgments.
Lank was supported under the ERC Advanced Grant 101095900-TriCatApp. Part of this work was completed while Lank visited the Basque Center for Applied Mathematics (BCAM) and thanks them for their hospitality. The authors thank Dylan Spence, Ana Cristina López-Martín, and Elías Guisado Villalgordo for discussions in an earlier version of the work.
@@ 2.1. Perfect & pseudocoherent complexes
We say $P\in D_{\operatorname{qc}}(X)$ is perfect if it is affine locally isomorphic to a bounded complex of locally free sheaves of finite rank. Denote by $\operatorname{Perf}(X)$ for the full subcategory of perfect complexes. Also, given an integer $N$, we say $E\in D_{\operatorname{qc}}(X)$ is $N$-pseudocoherent if locally there is a morphism $P\to E$ from some $P\in\operatorname{Perf}(X)$ such that the induced morphism on cohomology sheaves $\mathcal{H}^{n}(P)\to\mathcal{H}^{n}(E)$ is an isomorphism for $n>N$ and is surjective for $n=N$. We say $E\in D_{\operatorname{qc}}(X)$ is pseudocoherent if it is $N$-pseudocoherent for every $N\in\mathbb{Z}$. In general, since $X$ is a scheme, $E\in D_{\operatorname{qc}}(X)$ is pseudocoherent if and only if locally on $X$ there is a quasi-isomorphism $P\to E$, with $P$ a bounded above complex of finite-rank free modules [GW23, Corollary 22.47]. As $X$ is Noetherian, $E\in D_{\operatorname{qc}}(X)$ is pseudocoherent if, and only if, $\mathcal{H}^{n}(E)$ is coherent for all $n$ and vanishes for $n\gg 0$ (see e.g. [Sta25, Tag 08E8]). The subcategory of pseudocoherent complexes of $D_{\operatorname{qc}}(X)$ coincides with $D^{-}_{\operatorname{coh}}(X)$ (i.e. the full subcategory of complexes with bounded above and coherent cohomology). Consider a morphism $f\colon Y\to X$ of Noetherian schemes. Then one has $\mathbf{L}f^{\ast}D^{-}_{\operatorname{coh}}(X)\subseteq D^{-}_{\operatorname{coh}}(Y)$ (see e.g. [Sta25, Tag 08H4]). We will use this characterization liberally.
While straightforward, the following is useful.
@@
Let $f\colon Y\to X$ be an affine faithfully flat morphism of Noetherian schemes. An object $E\in D_{\operatorname{qc}}(X)$ belongs to $D^{-}_{\operatorname{coh}}(X)$ (resp. $\operatorname{Perf}(X)$) if, and only if, $\mathbf{L}f^{\ast}E\in D^{-}_{\operatorname{coh}}(Y)$ (resp. $\mathbf{L}f^{\ast}E\in\operatorname{Perf}(Y)$).
@@
This is a special case of [GW23, Proposition 22.52], but we spell it out for convenience. The problem is affine local. So, we can reduce to the affine case. Then the claim follows from [Sta25, Tag 068R] (resp. [GW23, Lemma 21.172]). ∎
@@ 2.2. Integral transforms
Now, we impose each $f_{i}$ be proper and flat. An object $K\in D^{-}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ is called relatively perfect over $Y_{1}$ if
@@
More generally, the definition of relatively perfect objects make sense for any proper morphism of Noetherian schemes $f\colon Y\to X$. Indeed, an object $K\in D^{-}_{\operatorname{coh}}(Y)$ is called $f$-perfect if $\mathbf{R}f_{\ast}(K\otimes^{\mathbf{L}}\operatorname{Perf}(Y))\subseteq\operatorname{Perf}(X)$. So, in our setting, $K\in D^{-}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ is relatively perfect over $Y_{1}$ precisely when it is $p_{1}$-perfect. This notion has appeared in the literature, but we record a lemma for convenience to avoid any potential confusion.
@@
    $E\otimes^{\mathbf{L}}\mathbf{L}f^{\ast}D^{b}_{\operatorname{coh}}(X)\subseteq D^{b}_{\operatorname{coh}}(Y)$.
@@
A useful observation from Lemma˜2.2 is that any object $E\in D^{-}_{\operatorname{coh}}(Y)$ being $f$-perfect implies it belongs to $D^{b}_{\operatorname{coh}}(Y)$.
@@ 2.3. Generation for triangulated categories
Let $\mathcal{T}$ be a triangulated category with shift functor $1]\colon\mathcal{T}\to\mathcal{T}$. Fix a subcategory $\mathcal{S}\subseteq\mathcal{T}$. Denote by $\operatorname{add}(\mathcal{S})$ the strictly full (that is, closed under isomorphisms) subcategory of $\mathcal{T}$ consisting of all direct summands of objects of the form $\oplus_{n\in\mathbb{Z}}S_{n}^{\oplus r_{n}}[n]$ where $S_{n}\in\mathcal{S}$ for all $n$ and finitely many $r_{n}$ are nonzero. Inductively, let
$\langle\mathcal{S}\rangle_{n}:=\begin{cases}\operatorname{add}(\varnothing)&\text{if }n=0,\\ \operatorname{add}(\mathcal{S})&\text{if }n=1,\\ \operatorname{add}(\{\operatorname{cone}\phi\mid\phi\in\operatorname{Hom}(\langle\mathcal{S}\rangle_{n-1},\langle\mathcal{S}\rangle_{1})\})&\text{if }n>1.\end{cases}$
Additionally, let $\langle\mathcal{S}\rangle:=\langle\mathcal{S}\rangle_{\infty}:=\cup_{n\geq 0}\langle\mathcal{S}\rangle_{n}$. It may be checked that $\langle\mathcal{S}\rangle$ is the smallest thick (that is, triangulated subcategory closed under direct summands) subcategory containing $\mathcal{S}$. In the case $\mathcal{S}$ consists of a single object $G$, we write the above as $\langle G\rangle_{n}$ and $\langle G\rangle$. An object $G\in\mathcal{T}$ is called a classical generator for $\mathcal{T}$ if $\langle G\rangle=\mathcal{T}$.
Now, we impose $\mathcal{T}$ be a compactly generated triangulated category. To clarify, we require that $\mathcal{T}$ admits all small coproducts. Denote the collection of compact objects of $\mathcal{T}$ by $\mathcal{T}^{c}$. We say $G\in\mathcal{T}^{c}$ is a compact generator for $\mathcal{T}$ if for any $E\in\mathcal{T}$, one has $\operatorname{Hom}(G,E[n])=0$ for all $n\in\mathbb{Z}$ if, and only if, $E=0$. A well-known fact is that classical generators for $\mathcal{T}^{c}$ coincide with being compact generators for $\mathcal{T}$ (see e.g. [Sta25, Tag 09SR]).
The following is an important case and is freely used without mention.
Example 2.4 ([BvdB03, Theorem 3.1.1]).
Let $X$ be a quasi-compact quasi-separated scheme. The collection of compact objects in $D_{\operatorname{qc}}(X)$ coincides with $\operatorname{Perf}(X)$. Moreover, $D_{\operatorname{qc}}(X)$ is compactly generated by a single object from $\operatorname{Perf}(X)$. See [BvdB03, Theorem 3.1.1] for details.
The following is known for some circles, but we add it for self-containedness (see e.g. [KL15, Lemma’s 2.10 & 2.12]).
@@
Remark 2.7.
Let $\Phi\colon D_{\operatorname{qc}}(Y_{1})\to D_{\operatorname{qc}}(Y_{2})$ be an exact functor where each $Y_{i}$ is Noetherian. Then the following are equivalent:
@@ 3. Preservation of adjointness
Consider ˜3.1 with $t$ a morphism locally of finite tor-dimension. Then $t_{1},t_{2},t^{\prime}$ are locally of finite tor-dimension. In particular, each face of Diagram˜3.1 is a tor-independent square (in the sense of [Lip09, Definition 3.10.2]).
@@
This first claim is affine local on $S$. Indeed, if needed, we can find an affine open cover of $S$ such that the base changes of $t$ along the associated open immersion are of finite tor-dimension. Then, for each such open immersion of a component in the affine cover, we have a similar diagram to Diagram˜3.1 (i.e. a ‘base change cube’). If we can show the desired properties on the morphisms arising in each diagram associated to the components of the affine open cover of $S$, then the desired claim follows. So, we may impose that $t$ be of finite tor-dimension.
By base change, we know that $g_{1},g_{2},f^{\prime}_{2},f^{\prime}_{1},g_{2}^{\prime},g_{1}^{\prime}$ are proper flat morphisms. Consider the fibered square in Diagram˜3.1 induced from the base change of $f_{1}$ along $t$. It follows by [GW23, Lemma 23.27] that $t_{1}$ is of finite tor-dimension. A similar argument shows $t_{2}$ is of finite tor-dimension as well. In fact, if we work with the fibered square in Diagram˜3.1 obtained by base change of $f_{2}^{\prime}$ along $t_{1}$, we see from similar reasoning that $t^{\prime}$ is locally of finite tor-dimension. This completes the proof for our first claim.
To check the second claim, it can be checked by one face at a time. Note that the same argument in what follows works whether $t$ has finite tor-dimension locally or absolutely because the conditions checked are at the level of stalks. The fibered square obtained by base change of $f_{1}$ along $t$ is tor-independent. Indeed, use [GW23, Exercise 21.32(iii)] with the constraint that $f_{1}$ is flat. Then [GW23, Exercise 21.33(v)], coupled with [Lip09, Definition 3.10.2(i) & Theorem 3.10.3], tells us the particular face is tor-independent. Now, we can proceed in a similar fashion with the remaining faces to finish the proof. ∎
@@
Consider ˜3.1 with $t$ an affine morphism. Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over $Y_{1}$. Then $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ restricts to an exact functor on $D^{b}_{\operatorname{coh}}$. Similarly, if $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ is relatively perfect over $Y_{2}$, then $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ restricts to an exact functor on $\operatorname{Perf}$.
@@
We only prove the case of $K$ being relatively perfect over $Y_{2}$ implying $\Phi_{\mathbf{L}(t^{\prime})K}$ restricts to an exact functor on $\operatorname{Perf}$ because a similar argument applies for $D^{b}_{\operatorname{coh}}$. As $\Phi_{K}(\operatorname{Perf}(Y_{1}))\subseteq\operatorname{Perf}(Y_{2})$, [DLM25, Proposition 3.2] implies
@@
Consider ˜3.1. Let $K\in D_{\operatorname{qc}}(Y_{1}\times_{S}Y_{2})$. Then there is a natural isomorphism $\beta^{K}\colon\mathbf{L}t_{2}^{\ast}\circ\Phi_{K}\to\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}\circ\mathbf{L}t_{1}^{\ast}$ on $D_{\operatorname{qc}}$.
@@
This follows from the string of functorial isomorphisms for each $E\in D_{\operatorname{qc}}(Y_{1}\times_{S}T)$:
$\mathbf{L}t_{2}^{\ast}\Phi_{K}(E)$ $\cong\mathbf{L}t_{2}^{\ast}\mathbf{R}(f^{\prime}_{1})_{\ast}(\mathbf{L}(f^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}K)$
$\cong\mathbf{R}(g^{\prime}_{1})_{\ast}\mathbf{L}(t^{\prime})^{\ast}(\mathbf{L}(f^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}K)$
$\cong\mathbf{R}(g^{\prime}_{1})_{\ast}(\mathbf{L}(t^{\prime})^{\ast}\mathbf{L}(f^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}\mathbf{L}(t^{\prime})^{\ast}K)$
$\cong\mathbf{R}(g^{\prime}_{1})_{\ast}(\mathbf{L}(g^{\prime}_{2})^{\ast}\mathbf{L}(t_{1}^{\ast}E\otimes^{\mathbf{L}}\mathbf{L}(t^{\prime})^{\ast}K))$
@@
Proposition 3.5.
Consider ˜3.1 where $t$ is an affine faithfully flat morphism. Let $K\in D^{-}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$. If $\mathbf{L}(t^{\prime})^{\ast}K$ is relatively perfect over $Y_{1}\times_{S}T$ (resp. $Y_{2}\times_{S}T$), then $K$ is relatively perfect over $Y_{1}$ (resp. $Y_{2}$).
@@
First, we show the case for $\mathbf{L}(t^{\prime})^{\ast}K$ being relatively perfect over $Y_{1}\times_{S}T$ implying $K$ is relatively perfect over $Y_{1}$. Choose $P_{i}\in\operatorname{Perf}(Y_{i})$ such that $\langle P_{i}\rangle=\operatorname{Perf}(Y_{i})$ for each $i$. From [BvdB03, Lemma 3.4.1], we know that $\langle\mathbf{L}f_{2}^{\ast}P_{1}\otimes^{\mathbf{L}}\mathbf{L}f_{1}^{\ast}P_{2}\rangle=\operatorname{Perf}(Y_{1}\times_{S}Y_{2})$. Then, by [Sta25, Tag 0BQT], it follows that
$\langle\mathbf{L}(t^{\prime})^{\ast}(\mathbf{L}f_{2}^{\ast}P_{1}\otimes^{\mathbf{L}}\mathbf{L}f_{1}^{\ast}P_{2})\rangle=\operatorname{Perf}(Y_{1}\times_{S}Y_{2}\times_{S}T).$
However, we have an isomorphism in Diagram˜3.1,
$\mathbf{L}(t^{\prime})^{\ast}(\mathbf{L}f_{2}^{\ast}P_{1}\otimes^{\mathbf{L}}\mathbf{L}f_{1}^{\ast}P_{2})$ $\cong\mathbf{L}(t^{\prime})^{\ast}\mathbf{L}f_{2}^{\ast}P_{1}\otimes^{\mathbf{L}}\mathbf{L}(t^{\prime})^{\ast}\mathbf{L}f_{1}^{\ast}P_{2}$
$\cong\mathbf{L}(g_{2}^{\prime})^{\ast}\mathbf{L}t_{1}^{\ast}P_{1}\otimes^{\mathbf{L}}\mathbf{L}(g_{1}^{\prime})^{\ast}\mathbf{L}t_{2}^{\ast}P_{2}.$
Now, we finish the proof for the claim. Let $E\in D^{b}_{\operatorname{coh}}(Y_{1})$. As $t^{\prime}$ is faithfully flat, it follows that $\mathbf{L}t_{1}^{\ast}E\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}T)$. By Lemma˜3.4, we know that $\mathbf{L}t_{2}^{\ast}\Phi_{K}(E)\cong\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}(\mathbf{L}t^{\ast}_{1}E)$. From Lemma˜2.1, the hypothesis implies $\Phi_{K}(E)\in D_{\operatorname{coh}}^{-}(Y_{2})$. So, it suffices to show that $\mathcal{H}^{j}(\Phi_{K}(E))=0$ for all but finitely many $j\in\mathbb{Z}$.
Our hypothesis ensures that $\Phi_{\mathbf{L}(t^{\prime})K}(\mathbf{L}t_{1}^{\ast}E)\in D^{b}_{\operatorname{coh}}(Y_{2}\times_{S}T)$. Then, as $t_{2}$ is flat, for each $j\in\mathbb{Z}$,

### v3

@@ Acknowledgments.
Lank was supported under the ERC Advanced Grant 101095900-TriCatApp. Part of this work was completed while Lank visited the Basque Center for Applied Mathematics (BCAM) and thanks them for their hospitality. Guisado Villalgordo was supported under the grant PRE2022-104796 by the Spanish Ministry of Science and Innovation. Manali Rahul is grateful to Max Planck Institute for Mathematics in Bonn for its hospitality and financial support. The authors thank Dylan Spence, Ana Cristina López-Martín, and Fernando Sancho de Salas for discussions.
@@ 2.1. Perfect & pseudocoherent complexes
We say $P\in D_{\operatorname{qc}}(X)$ is perfect if it is locally isomorphic to a bounded complex of locally free sheaves of finite rank. Denote by $\operatorname{Perf}(X)$ for the full subcategory of perfect complexes. Also, given an integer $N$, we say $E\in D_{\operatorname{qc}}(X)$ is $N$-pseudocoherent if locally there is a morphism $P\to E$ from some $P\in\operatorname{Perf}(X)$ such that the induced morphism on cohomology sheaves $\mathcal{H}^{n}(P)\to\mathcal{H}^{n}(E)$ is an isomorphism for $n>N$ and is surjective for $n=N$. We say $E\in D_{\operatorname{qc}}(X)$ is pseudocoherent if it is $N$-pseudocoherent for every $N\in\mathbb{Z}$.
In general, $E\in D_{\operatorname{qc}}(X)$ is pseudocoherent if, and only if, locally on $X$ there is a quasi-isomorphism $P\to E$, with $P$ a bounded above complex of finite-rank free modules [GW23, Corollary 22.47]. As $X$ is Noetherian, $E\in D_{\operatorname{qc}}(X)$ is pseudocoherent if, and only if, $\mathcal{H}^{n}(E)$ is coherent for all $n$ and vanishes for $n\gg 0$ (see e.g. [STA25, Tag 08E8]). The subcategory of pseudocoherent complexes of $D_{\operatorname{qc}}(X)$ coincides with $D^{-}_{\operatorname{coh}}(X)$ (i.e. the full subcategory of complexes with bounded above and coherent cohomology).
Consider a morphism $f\colon Y\to X$ of Noetherian schemes. Then one has $\mathbf{L}f^{\ast}D^{-}_{\operatorname{coh}}(X)\subseteq D^{-}_{\operatorname{coh}}(Y)$ (see e.g. [STA25, Tag 08H4]). We record a few straightforward results.
@@
Let $f\colon Y\to X$ be a faithfully flat morphism of Noetherian schemes. An object $E\in D_{\operatorname{qc}}(X)$ belongs to $D^{-}_{\operatorname{coh}}(X)$ (resp. $\operatorname{Perf}(X)$) if, and only if, $\mathbf{L}f^{\ast}E\in D^{-}_{\operatorname{coh}}(Y)$ (resp. $\mathbf{L}f^{\ast}E\in\operatorname{Perf}(Y)$).
@@
This is a special case of [GW23, Proposition 22.52]. ∎
Lemma 2.2.
Let $f\colon W\to Y$ be a morphism of schemes. Choose $E,G\in D(Y)$. Assume one of the following holds:
  1. (1)
    $E$ is perfect
  2. (2)
    $E$ is pseudocoherent and both $G,\mathbf{L}f^{\ast}G$ are locally bounded below.
Then the natural morphism
$\mathbf{L}f^{\ast}\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(E,G)\to\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}f^{\ast}E,\mathbf{L}f^{\ast}G)$
is an isomorphism.
Proof.
On one hand, (1) is a special case of [GW23, Proposition 22.70]. On the other hand, (2) is argued exactly as in the case of loc. cit.  where ‘$f$ has finite tor-dimension’ can be replaced with our hypothesis that $\mathbf{L}f^{\ast}G$ is locally bounded below. Indeed, this is the only part in the proof of loc. cit. which requires the assumption of finite tor-dimension. ∎
Next, we give a characterization of perfectness. To do so, we need notation. Given a scheme $X$ and a point $p\in X$, denote by $\sigma_{p}\colon\operatorname{Spec}(\mathcal{O}_{X,p})\to X$ and $i_{p}\colon\operatorname{Spec}(\kappa(p))\to X$ for the natural morphisms. Note that the set theoretic image of $\sigma_{p}$ is the intersection of all open subschemes $U$ of $X$ which contain $p$.
Lemma 2.3.
Let $X$ be a scheme, $E\in D_{\operatorname{qc}}^{b}(X)$ be pseudocoherent, and $p\in X$. Then the following are equivalent:
  1. (1)
    There is an open immersion $j\colon U\to X$ such that $p\in U$ and $\mathbf{L}j^{\ast}E$ is perfect
  2. (2)
    $\mathbf{L}\sigma_{p}^{\ast}E$ is perfect
  3. (3)
    $\mathbf{L}i_{p}^{\ast}E$ is bounded.
Proof.
[…]
$\operatorname{Spec}(\kappa(p))\to\operatorname{Spec}(\mathcal{O}_{X,p})\to U\xrightarrow{j}X.$
Indeed, one uses the fact that derived pullback preserves perfect complexes (see e.g. [STA25, Tag 09UA]).
[…]
The following is an extension of [AJS23, Theorem 2.3] to closed points.
Proposition 2.4.
Let $X$ be a quasi-separated quasi-compact scheme and $E\in D^{b}_{\operatorname{qc}}(X)$ be pseudocoherent. Then the following conditions are equivalent:
  1. (1)
    $E$ is perfect
  2. (2)
    $\mathbf{L}i_{p}^{\ast}E$ is bounded for every $p\in X$
  3. (3)
    $\mathbf{L}i_{p}^{\ast}E$ is bounded for every $p\in X$ which is closed
  4. (4)
    $\mathbf{L}\sigma_{p}^{\ast}E$ is perfect for every $p\in X$
  5. (5)
    $\mathbf{L}\sigma_{p}^{\ast}E$ is perfect for every $p\in X$ which is closed.
Proof.
[…]
@@ 2.2. Integral transforms
Now, suppose each $f_{i}$ be proper and flat. An object $K\in D^{-}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ is called relatively perfect over $Y_{1}$ if
@@
More generally, the definition of relatively perfect objects make sense for any proper morphism of Noetherian schemes $f\colon Y\to X$. Indeed, an object $K\in D^{-}_{\operatorname{coh}}(Y)$ is called $f$-perfect if $\mathbf{R}f_{\ast}(K\otimes^{\mathbf{L}}\operatorname{Perf}(Y))\subseteq\operatorname{Perf}(X)$. So, in our setting, $K\in D^{-}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ is relatively perfect over $Y_{1}$ precisely when it is $p_{1}$-perfect. This notion has appeared in the literature but we record a lemma for convenience to avoid any potential confusion.
@@
    $E\otimes^{\mathbf{L}}\mathbf{L}f^{\ast}D^{b}_{\operatorname{coh}}(X)\subseteq D^{b}_{\operatorname{coh}}(Y)$
@@
A useful observation from Lemma˜2.5 is that any object $E\in D^{-}_{\operatorname{coh}}(Y)$ being $f$-perfect implies it belongs to $D^{b}_{\operatorname{coh}}(Y)$.
@@ 2.3. Generation for triangulated categories
Let $\mathcal{T}$ be a triangulated category with shift functor $1]\colon\mathcal{T}\to\mathcal{T}$. Fix a subcategory $\mathcal{S}\subseteq\mathcal{T}$. Denote by $\operatorname{add}(\mathcal{S})$ for the strictly full (that is, closed under isomorphisms) subcategory of $\mathcal{T}$ consisting of all direct summands of objects of the form $\oplus_{n\in\mathbb{Z}}S_{n}^{\oplus r_{n}}[n]$ where $S_{n}\in\mathcal{S}$ for all $n$ and finitely many $r_{n}$ are nonzero. Inductively, let
$\langle\mathcal{S}\rangle_{n}:=\begin{cases}\operatorname{add}(\varnothing)&\text{if }n=0\\ \operatorname{add}(\mathcal{S})&\text{if }n=1\\ \operatorname{add}(\{\operatorname{cone}\phi\mid\phi\in\operatorname{Hom}(\langle\mathcal{S}\rangle_{n-1},\langle\mathcal{S}\rangle_{1})\})&\text{if }n>1.\end{cases}$
Additionally, let $\langle\mathcal{S}\rangle:=\cup_{n\geq 0}\langle\mathcal{S}\rangle_{n}$. It may be checked that $\langle\mathcal{S}\rangle$ is the smallest thick (that is, triangulated subcategory closed under direct summands) subcategory containing $\mathcal{S}$. In the case $\mathcal{S}$ consists of a single object $G$, we write the above as $\langle G\rangle_{n}$ and $\langle G\rangle$. An object $G\in\mathcal{T}$ is called a classical generator for $\mathcal{T}$ if $\langle G\rangle=\mathcal{T}$.
Suppose that $\mathcal{T}$ is compactly generated triangulated category (to clarify, we require that $\mathcal{T}$ admits all small coproducts in the definition of being compactly generated). Denote the collection of compact objects of $\mathcal{T}$ by $\mathcal{T}^{c}$. We say $G\in\mathcal{T}^{c}$ is a compact generator for $\mathcal{T}$ if for any $E\in\mathcal{T}$, one has $\operatorname{Hom}(G,E[n])=0$ for all $n\in\mathbb{Z}$ if, and only if, $E=0$. A well-known fact is that classical generators for $\mathcal{T}^{c}$ coincide with compact generators for $\mathcal{T}$ (see e.g. [STA25, Tag 09SR]).
The following is an important case.
Example 2.7 ([Bv03, Theorem 3.1.1]).
Let $X$ be a quasi-compact quasi-separated scheme. The subcategory of compact objects in $D_{\operatorname{qc}}(X)$ coincides with $\operatorname{Perf}(X)$. Moreover, $D_{\operatorname{qc}}(X)$ is compactly generated by a single object from $\operatorname{Perf}(X)$, see [Bv03, Theorem 3.1.1] for details.
We add the following for self-containedness (see e.g. [KL15, Lemma’s 2.10 & 2.12]).
@@
Lemma 2.10.
Let $Y_{1}$ and $Y_{2}$ be Noetherian schemes. Suppose $\Phi\colon D_{\operatorname{qc}}(Y_{1})\to D_{\operatorname{qc}}(Y_{2})$ is an exact functor which preserves small coproducts. Then the following are equivalent:
@@ 3. Preservation of adjointness
Consider ˜3.1. Then each face of Diagram˜3.1 is a tor-independent square (in the sense of [LIP09, Definition 3.10.2]).
@@
As each $f_{i}$, $f^{\prime}_{i}$, $g_{i}$, and $g^{\prime}_{i}$ are flat, the desired claim follows from [GW23, Remark 22.94]. ∎
@@
Consider ˜3.1 with $t$ an affine morphism. If $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ is relatively perfect over $Y_{1}$ (resp. over $Y_{2}$), then $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ restricts to an exact functor on $D^{b}_{\operatorname{coh}}$ (resp. on $\operatorname{Perf}$).
@@
We only prove the case of $K$ being relatively perfect over $Y_{2}$ because a similar argument applies for $D^{b}_{\operatorname{coh}}$. As $\Phi_{K}(\operatorname{Perf}(Y_{1}))\subseteq\operatorname{Perf}(Y_{2})$, [DLM25, Proposition 3.2] implies
@@
Consider ˜3.1. Let $K\in D_{\operatorname{qc}}(Y_{1}\times_{S}Y_{2})$. Then on $D_{\operatorname{qc}}$ there is a natural isomorphism
$\beta^{K}\colon\mathbf{L}t_{2}^{\ast}\circ\Phi_{K}\to\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}\circ\mathbf{L}t_{1}^{\ast}.$
@@
This follows from the string of isomorphisms for each $E\in D_{\operatorname{qc}}(Y_{1}\times_{S}T)$:
$\mathbf{L}t_{2}^{\ast}\Phi_{K}(E)$ $=\mathbf{L}t_{2}^{\ast}\mathbf{R}(f^{\prime}_{1})_{\ast}(\mathbf{L}(f^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}K)$
$\cong\mathbf{R}(g^{\prime}_{1})_{\ast}\mathbf{L}(t^{\prime})^{\ast}(\mathbf{L}(f^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}K)$ (flat base change)
$\cong\mathbf{R}(g^{\prime}_{1})_{\ast}(\mathbf{L}(t^{\prime})^{\ast}\mathbf{L}(f^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}\mathbf{L}(t^{\prime})^{\ast}K)$ (monoidality of $\mathbf{L}(-)^{\ast}$)
$\cong\mathbf{R}(g^{\prime}_{1})_{\ast}(\mathbf{L}(g^{\prime}_{2})^{\ast}\mathbf{L}(t_{1})^{\ast}E\otimes^{\mathbf{L}}\mathbf{L}(t^{\prime})^{\ast}K)$ (pseudofunctoriality of $\mathbf{L}(-)^{\ast}$)
@@
Lemma 3.5.
Consider ˜3.1. Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$.
  1. (1)
    If $\mathbf{L}(t^{\prime})^{\ast}K$ is relatively perfect over $Y_{1}\times_{S}T$ (resp. $Y_{2}\times_{S}T$) where $t$ is affine and faithfully flat, then $K$ is relatively perfect over $Y_{1}$ (resp. $Y_{2}$).
  2. (2)
    If $\mathbf{L}(t^{\prime})^{\ast}K$ is relatively perfect over $Y_{1}\times_{S}\operatorname{Spec}(k)$ (resp. $Y_{2}\times_{S}\operatorname{Spec}(k)$) for all morphisms $\operatorname{Spec}(k)\to S$ from a field, then $K$ is relatively perfect over $Y_{1}$ (resp. $Y_{2}$).
  3. (3)
    If $\mathbf{L}(t^{\prime})^{\ast}K$ is relatively perfect over $Y_{1}\times_{S}\operatorname{Spec}(\kappa(p))$ (resp. $Y_{2}\times_{S}\operatorname{Spec}(\kappa(p))$) for all closed immersions $\operatorname{Spec}(\kappa(p))\to S$ associated to a closed point $p\in S$, then $K$ is relatively perfect over $Y_{1}$ (resp. $Y_{2}$).
@@
We only show the case for $\mathbf{L}(t^{\prime})^{\ast}K$ being relatively perfect over $Y_{1}\times_{S}T$. To show the other case, one can argue like below and use Lemma˜2.1.
To start, we prove the first claim. Let $E\in D^{b}_{\operatorname{coh}}(Y_{1})$. As $t^{\prime}$ is faithfully flat, it follows that $\mathbf{L}t_{1}^{\ast}E\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}T)$. By Lemma˜3.4, we know that $\mathbf{L}t_{2}^{\ast}\Phi_{K}(E)\cong\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}(\mathbf{L}t^{\ast}_{1}E)$. From Lemma˜2.1, the hypothesis implies $\Phi_{K}(E)\in D_{\operatorname{coh}}^{-}(Y_{2})$. So, it suffices to show that $\mathcal{H}^{j}(\Phi_{K}(E))=0$ for all but finitely many $j\in\mathbb{Z}$. Our hypothesis ensures that $\Phi_{\mathbf{L}(t^{\prime})K}(\mathbf{L}t_{1}^{\ast}E)\in D^{b}_{\operatorname{coh}}(Y_{2}\times_{S}T)$. Then, as $t_{2}$ is flat, for each $j\in\mathbb{Z}$,